# Ground States of ${\text{Class}\;\mathcal{S}}$ Theory on ADE Singularities and dual Chern-Simons theory

### v1

@@ B.1 Contribution from vector multiplet $I^{vec}$
$(F^{0}+F^{+})(t^{2}-t^{4}+2t^{6})=3F^{0}(t^{2}-t^{4}+2t^{6}).$ (68)
Now let us consider the $j_{1}=\pm 1/2$ contribution. A generic operator now takes the form:
$\sum c^{p\mu n_{1}n_{2}}\partial^{n_{1}}_{++}\partial^{n_{2}}_{-+}\lambda_{\mu p},$
where recall that $\mu$ index is for spin $j_{1}$ and $p$ is SU$(2)$ adjoint index. The constraints are
$\sum e^{\pi i/2(n_{1}-n_{2}+\mu)}c^{p\mu n_{1}n_{2}}\partial^{n_{1}}_{++}\partial^{n_{2}}_{-+}\lambda_{\mu p}$ $=\sum(-1)^{p}c^{p\mu n_{1}n_{2}}\partial^{n_{1}}_{++}\partial^{n_{2}}_{-+}\lambda_{\mu p},$
$\sum(-1)^{(\mu-1)/2}e^{\pi in_{2}}c^{p\mu n_{1}n_{2}}\partial^{n_{2}}_{++}\partial^{n_{1}}_{-+}\lambda_{-\mu p}$ $=-\sum c^{p\mu n_{1}n_{2}}\partial^{n_{1}}_{++}\partial^{n_{2}}_{-+}\lambda_{\mu,-p}.$
@@
$e^{\pi i/2(n_{1}-n_{2}+\mu)}c^{p\mu n_{1}n_{2}}$ $=(-1)^{p}c^{p\mu n_{1}n_{2}},$
$(-1)^{(-\mu-1)/2}e^{\pi in_{1}}c^{-p,-\mu,n_{2}n_{1}}$ $=-c^{p\mu n_{1}n_{2}}.$
There are now three cases to consider:
Case 1: $p=0,\mu=1$. For this case, the first constraint implies
$n_{1}-n_{2}=3\mod{4},$
and the second constraint shows that the case $p=0,\mu=-1$ is taken care of if we sum over all such $n_{1}$ $n_{2}$ pairs. The sum is
$F^{01}=\frac{t^{3}+t^{9}}{(1-t^{6})(1-t^{12})}.$ (69)
Case 2: $p=1,\mu=1$. For this case, the first constraint implies
$n_{1}-n_{2}=1\mod{4},$
and the second constraint shows that the case $p=-1,\mu=-1$ is taken care of if we sum over all such $n_{1}$ $n_{2}$ pairs. The sum is
$F^{11}=\frac{t^{3}+t^{9}}{(1-t^{6})(1-t^{12})}=F^{01}.$ (70)
Case 3: $p=1,\mu=-1$. For this case, the first constraint implies
$n_{1}-n_{2}=3\mod{4},$
and the second constraint shows that the case $p=-1,\mu=1$ is taken care of if we sum over all such $n_{1}$ $n_{2}$ pairs. The sum is
$F^{1,-1}=\frac{t^{3}+t^{9}}{(1-t^{6})(1-t^{12})}=F^{01}.$ (71)
Adding this up, we find that the $j_{1}=\pm 1/2$ contribution to the single letter vector multiplet index is
$-(F^{01}+F^{11}+F^{1,-1})t^{3}=-3F^{01}t^{3}.$ (72)
@@
$I^{vec}_{a}$ $=\frac{3(1-t^{6}+t^{12})(t^{2}-t^{4}+2t^{6})-3(t^{3}+t^{9})t^{3}}{(1-t^{6})(1-t^{12})},$ (73)
$I^{vec}_{5}$ $=3\frac{-t^{6}+t^{8}-t^{10}+t^{12}}{(1-t^{6})(1-t^{12})},$ (74)
with $a=\{1,2,3,4\}$.
@@
B.2 Contribution from half-hypermultiplet $I^{trifund}$
We go over the explicit calculation of the single letter index associated to the trifundamental field $I^{trifund}_{abc}$ where $a,b,c\in\{1,2,3,4\}$. According to equation (58), there are only 4 independent combinations that give unique constraints: $(111)$, $(112)$, $(113)$, $(114)$ (for example, $(122)$ would be equivalent to $(111)$ in terms of constraints on the trifundamental fields), so we only have four independant indices to compute. Note that all component fields in the half-hypermultiplet have $j_{1}=0$, so we need not worry about the $j_{1}=\pm 1/2$ contributions.
In the case of $I^{trifund}_{111}$, a generic operator that arises from the state-operator correspondence now has the form:
$\sum c^{n_{1}n_{2}}\partial^{n_{1}}_{++}\partial^{n_{2}}_{-+}\Phi,$
where $\Phi$ denotes trifundamental field. For clarity, we suppress the SU$(2)$ trifundamental indices on the field $\Phi$, which otherwise would look like as in equation (58). The $r$ and the $s$ constraints give
$c^{n_{1}n_{2}}$ $=e^{i\pi/2(n_{1}-n_{2})}c^{n_{1}n_{2}},$
$c^{n_{1}n_{2}}$ $=e^{i\pi n_{1}}c^{n_{2}n_{1}}.$
@@ B.2 Contribution from half-hypermultiplet $I^{trifund}$
$n_{1}-n_{2}=0\mod{4},$
and the second constraint instructs us to sum over $n_{1}\geq n_{2}$ only, but for $n_{1}=n_{2}$ we need $n_{1}=0\mod{2}$. Overall, summing over the Boltzmann weights, we have
$F_{111}=8\frac{1-t^{6}+t^{12}}{(1-t^{6})(1-t^{12})},$
where the prefactor 8 comes from $2^{3}$ choices of the SU$(2)$ trifundamental indices. Therefore,
$I^{trifund}_{111}=F_{111}(t^{2}-t^{4})=8\frac{(1-t^{6}+t^{12})(t^{2}-t^{4})}{(1-t^{6})(1-t^{12})}.$ (75)
In the case of $I^{trifund}_{112}$, the constraints are
$e^{i\pi/2(n_{1}-n_{2})}c^{n_{1}n_{2}}$ $=c^{n_{1}n_{2}},$
$e^{i\pi n_{1}}c^{n_{2}n_{1}}$ $=-c^{n_{1}n_{2}},$
which gives
$F_{112}=8\frac{t^{6}}{(1-t^{6})(1-t^{12})},$
@@
$I^{trifund}_{112}=F_{112}(t^{2}-t^{4})=8\frac{t^{6}(t^{2}-t^{4})}{(1-t^{6})(1-t^{12})}.$ (76)
In the case of $I^{trifund}_{113}$, the constraints are
$e^{i\pi/2(n_{1}-n_{2})}c^{n_{1}n_{2}}$ $=-c^{n_{1}n_{2}},$
$e^{i\pi n_{1}}c^{n_{2}n_{1}}$ $=c^{n_{1}n_{2}}.$
@@
$n_{1}-n_{2}=2\mod{4},$
and the second one tells us to sum over all $n_{1},n_{2}$ pairs. This gives
$F_{113}=8\frac{t^{6}}{(1-t^{6})(1-t^{12})}=F_{112},$
and
$I^{trifund}_{113}=8F_{113}(t^{2}-t^{4})=8\frac{t^{6}(t^{2}-t^{4})}{(1-t^{6})(1-t^{12})}.$ (77)
In the case of $I^{trifund}_{114}$, the constraints are
$e^{i\pi/2(n_{1}-n_{2})}c^{n_{1}n_{2}}$ $=-c^{n_{1}n_{2}},$
$e^{i\pi n_{1})}c^{n_{2}n_{1}}$ $=-c^{n_{1}n_{2}}.$
Comparing the (114) constraints with the (113) constraints, we see that they have the same solutions, so
$I^{trifund}_{114}=I^{trifund}_{113}=8\frac{t^{6}(t^{2}-t^{4})}{(1-t^{6})(1-t^{12})}.$ (78)
@@
Now, we study the index when one/two/three of the holonomies in the triplet are the 5th holonomy. First, consider the case where there is only one non-abelian holonomy. Similar to the argument we had previously, there are only 4 independent cases to consider: (115), (125), (135), (145).
In the case of $I^{trifund}_{115}$, the SU$(2)$ index of the third component of the trifundamental field matters, because the nonabelian holonomy does not act diagonally anymore. This consideration leads to the most general operator as in
$\sum c^{\mu n_{1}n_{2}}\partial^{n_{1}}_{++}\partial^{n_{2}}_{-+}\Phi_{\mu},$
where $\mu=\pm$ is the third component of the trifundamental index. For clarity, we keep first two components hidden. The constraints are
$\sum e^{i\pi/2(n_{1}-n_{2})}c^{\mu n_{1}n_{2}}\partial^{n_{1}}_{++}\partial^{n_{2}}_{-+}\Phi_{\mu}$ $=\sum e^{\mu\pi i/2}c^{\mu n_{1}n_{2}}\partial^{n_{1}}_{++}\partial^{n_{2}}_{-+}\Phi_{\mu},$
$\sum e^{i\pi n_{2}}c^{\mu n_{1}n_{2}}\partial^{n_{2}}_{++}\partial^{n_{1}}_{-+}\Phi_{\mu}$ $=\sum(-1)^{(\mu-1)/2}c^{\mu n_{1}n_{2}}\partial^{n_{1}}_{++}\partial^{n_{2}}_{-+}\Phi_{-\mu},$
@@
$e^{i\pi/2(n_{1}-n_{2})}c^{\mu n_{1}n_{2}}$ $=e^{\mu\pi i/2}c^{\mu n_{1}n_{2}},$
$e^{i\pi n_{1}}c^{-\mu,n_{2}n_{1}}$ $=(-1)^{(\mu-1)/2}c^{\mu n_{1}n_{2}}.$
For $\mu=1$, the first constraint implies
$n_{1}-n_{2}=1\mod{4},$
and the second constraint tells us to sum over all $n_{1},n_{2}$ pairs, which would take care of the case $\mu=-1$. Therefore, summing over the Boltzmann weights, we get
$F_{115}=4\frac{t^{3}+t^{9}}{(1-t^{6})(1-t^{12})},$
and
$I^{trifund}_{115}=F_{115}(t^{2}-t^{4})=4\frac{(t^{3}+t^{9})(t^{2}-t^{4})}{(1-t^{6})(1-t^{12})},$ (79)
where the $4$ comes from $2^{2}$ choices of the first two SU$(2)$ trifundamental indices. The factor $t^{2}-t^{4}$ comes from the contribution of the half-hypermultiplet base fields $q^{1}$ and $\bar{\psi}_{\dot{+}}$.
In the case of $I^{trifund}_{125}$, the constraints are
$\sum e^{i\pi/2(n_{1}-n_{2})}c^{\mu n_{1}n_{2}}\partial^{n_{1}}_{++}\partial^{n_{2}}_{-+}\Phi_{\mu}$ $=\sum e^{\mu\pi i/2}c^{\mu n_{1}n_{2}}\partial^{n_{1}}_{++}\partial^{n_{2}}_{-+}\Phi_{\mu},$
$\sum e^{i\pi n_{2}}c^{\mu n_{1}n_{2}}\partial^{n_{2}}_{++}\partial^{n_{1}}_{-+}\Phi_{\mu}$ $=-\sum(-1)^{(\mu-1)/2}c^{\mu n_{1}n_{2}}\partial^{n_{1}}_{++}\partial^{n_{2}}_{-+}\Phi_{-\mu}.$
The only difference from the previous $115$ case is the minus sign for the $s$ condition, so
$I^{trifund}_{125}=I^{trifund}_{115}=4\frac{(t^{3}+t^{9})(t^{2}-t^{4})}{(1-t^{6})(1-t^{12})}.$ (80)
In the case of $I^{trifund}_{135}$, the constraints are
$\sum e^{i\pi/2(n_{1}-n_{2})}c^{\mu n_{1}n_{2}}\partial^{n_{1}}_{++}\partial^{n_{2}}_{-+}\Phi_{\mu}$ $=-\sum e^{\mu\pi i/2}c^{\mu n_{1}n_{2}}\partial^{n_{1}}_{++}\partial^{n_{2}}_{-+}\Phi_{\mu},$
$\sum e^{i\pi n_{2}}c^{\mu n_{1}n_{2}}\partial^{n_{2}}_{++}\partial^{n_{1}}_{-+}\Phi_{\mu}$ $=\sum(-1)^{(\mu-1)/2}c^{\mu n_{1}n_{2}}\partial^{n_{1}}_{++}\partial^{n_{2}}_{-+}\Phi_{-\mu}.$
@@
$I^{trifund}_{135}=I^{trifund}_{115}=4\frac{(t^{3}+t^{9})(t^{2}-t^{4})}{(1-t^{6})(1-t^{12})}.$ (81)
@@
$I^{trifund}_{145}=I^{trifund}_{115}=4\frac{(t^{3}+t^{9})(t^{2}-t^{4})}{(1-t^{6})(1-t^{12})}.$ (82)
@@
This completes the computation for the case when only one of the three holonomies is nonabelian. We now consider the four cases where two of the holonomies are nonabelian. Those are: (155), (255), (355), and (455).
@@
In the case of $I^{trifund}_{155}$, the second and the third SU$(2)$ trifundamental indices matter, since the holonomies will act nondiagonally on them separately. So the general operator combination is
$\sum c^{\mu\nu n_{1}n_{2}}\partial^{n_{1}}_{++}\partial^{n_{2}}_{-+}\Phi_{\mu\nu},$
where we do not show the first SU$(2)$ trifundamental index. The constraints are
$\sum e^{\pi i/2(n_{1}-n_{2})}c^{\mu\nu n_{1}n_{2}}\partial^{n_{1}}_{++}\partial^{n_{2}}_{-+}\Phi_{\mu\nu}$ $=\sum e^{\pi(\mu+\nu)i/2}c^{\mu\nu n_{1}n_{2}}\partial^{n_{1}}_{++}\partial^{n_{2}}_{-+}\Phi_{\mu\nu},$
$\sum e^{\pi in_{2}}c^{\mu\nu n_{1}n_{2}}\partial^{n_{2}}_{++}\partial^{n_{1}}_{-+}\Phi_{\mu\nu}$ $=\sum(-1)^{(\mu+\nu)/2-1}c^{\mu\nu n_{1}n_{2}}\partial^{n_{1}}_{++}\partial^{n_{2}}_{-+}\Phi_{-\mu,-\nu},$
from which we have
$e^{\pi i/2(n_{1}-n_{2})}c^{\mu\nu n_{1}n_{2}}$ $=e^{\pi(\mu+\nu)i/2}c^{\mu\nu n_{1}n_{2}},$
$e^{\pi in_{1}}c^{-\mu,-\nu,n_{2}n_{1}}$ $=(-1)^{(\mu+\nu)/2-1}c^{\mu\nu n_{1}n_{2}}.$
Case 1: $\mu=\nu=1$. The first constraint implies
$n_{1}-n_{2}=2\mod{4},$
and the second constraint says that summing over all such pairs will help us take care of the case $\mu=\nu=-1$ case. The Boltzmann weight sum is
$F_{11}=2\frac{2t^{6}}{(1-t^{6})(1-t^{12})},$

### v2

@@ B.1 Contribution of a single vector multiplet $I^{vec}$
$F^{0}+F^{+}=\frac{3t^{6}(t^{2}-t^{4}+2t^{6})}{(1-t^{6})(1-t^{12})}.$ (71)
Now let us consider the $j_{1}=\pm 1/2$ contribution. A generic operator of this kind takes the form
$\sum c^{p\mu n_{1}n_{2}}\partial^{n_{1}}_{+\dot{+}}\partial^{n_{2}}_{-\dot{+}}\lambda_{\mu p},$
where we recall that $\mu=\pm 1$ labels the $j_{1}=\mu/2$ and $p=-1,0,1$ labels the $\mathfrak{su}(2)$ Lie algebra component. The constraints are
$\sum e^{i\pi(n_{1}-n_{2}+\mu)/2}c^{p\mu n_{1}n_{2}}\partial^{n_{1}}_{+\dot{+}}\partial^{n_{2}}_{-\dot{+}}\lambda_{\mu p}$ $=\sum(-1)^{p}c^{p\mu n_{1}n_{2}}\partial^{n_{1}}_{+\dot{+}}\partial^{n_{2}}_{-\dot{+}}\lambda_{\mu p},$
$\sum(-1)^{n_{2}+(\mu-1)/2}c^{p\mu n_{1}n_{2}}\partial^{n_{2}}_{+\dot{+}}\partial^{n_{1}}_{-\dot{+}}\lambda_{-\mu p}$ $=-\sum c^{p\mu n_{1}n_{2}}\partial^{n_{1}}_{+\dot{+}}\partial^{n_{2}}_{-\dot{+}}\lambda_{\mu,-p}.$
@@
$e^{i\pi(n_{1}-n_{2}+\mu)/2}c^{p\mu n_{1}n_{2}}=(-1)^{p}c^{p\mu n_{1}n_{2}},\qquad(-1)^{n_{1}-(\mu+1)/2}c^{-p,-\mu,n_{2}n_{1}}=-c^{p\mu n_{1}n_{2}}.$ (72)
The second equation of (72) expresses $c^{(-1)\mu n_{1}n_{2}}$ in terms of $c^{1(-\mu)n_{2}n_{1}}$, and so we only need to count independent coefficients $c^{p\mu n_{1}n_{2}}$ for $p=0,1$. We separate the calculation into four cases.
  • For $p=0,\mu=1$, the first constraint of (72) requires
    $n_{1}-n_{2}\equiv 3\pmod{4}.$
    The generating function for the number of such pairs $(n_{1},n_{2})$, each pair counted as $t^{3(n_{1}+n_{2})}$, is
    $\frac{t^{3}+t^{9}}{(1-t^{6})(1-t^{12})},$
    and the contribution to the index is
    $F^{01}=-\frac{t^{3}(t^{3}+t^{9})}{(1-t^{6})(1-t^{12})}.$ (73)
    where the prefactor $(-t^{3})$ corresponds to the quantum numbers of $\lambda^{1}_{\mu}$, as listed in Table 3.
  • For $p=0,\mu=-1$, since the second constraint of (72) determines $c^{0(-1)n_{1}n_{2}}$ in terms of $c^{01n_{2}n_{1}}$, there are no additional contributions to the index that have not been accounted for by the previous case, $p=0,\mu=1$.
  • For $p=1,\mu=1$, the first constraint of (72) implies
    $n_{1}-n_{2}\equiv 1\pmod{4}.$
    The sum is therefore
    $F^{11}=-\frac{t^{3}(t^{3}+t^{9})}{(1-t^{6})(1-t^{12})}=F^{01}.$ (74)
  • For $p=1,\mu=-1$, the first constraint of (72) implies
    $n_{1}-n_{2}\equiv 3\pmod{4}.$
    The sum is therefore
    $F^{1,-1}=-\frac{t^{3}(t^{3}+t^{9})}{(1-t^{6})(1-t^{12})}=F^{01}.$ (75)
Adding all four contributions together, we find that the $j_{1}=\pm 1/2$ contribution to the single letter vector multiplet index is
$F^{01}+F^{11}+F^{1,-1}=-\frac{3t^{3}(t^{3}+t^{9})}{(1-t^{6})(1-t^{12})}.$ (76)
@@
$I^{vec}_{a}$ $=\frac{3t^{2}(1-t^{6}+t^{12})(1-t^{2}+2t^{4})-3(t^{3}+t^{9})t^{3}}{(1-t^{6})(1-t^{12})},\qquad(a=1,\dots,4),$ (77)
$I^{vec}_{5}$ $=-3\frac{t^{6}(1+t^{4})(1-t^{2})}{(1-t^{6})(1-t^{12})}.$ (78)
@@
B.2 The contribution $I^{\text{(trifund)}}$ of the half-hypermultiplet
We go over the explicit calculation of the single letter index $I^{\text{(trifund)}}_{abc}$ associated to the trifundamental field, where $a,b,c\in\{1,2,3,4\}$. According to equation (58) and Table 4, there are only 4 independent combinations of $(abc)$ that give unique constraints, which we can take to be $(111)$, $(112)$, $(113)$, $(114)$, for example. The holonomy triplet $(122)$, for instance, is equivalent to $(111)$ in terms of the implied constraints on the trifundamental fields, since the factors $g_{a}(r)g_{b}(r)g_{c}(r)$ and $g_{a}(s)g_{b}(s)g_{c}(s)$ on the left hand side of (58) are the same for both $(a,b,c)=(1,1,1)$ and $(a,b,c)=(1,2,2)$. Therefore, we only have 4 independent indices to compute. We also note that all component fields in the half-hypermultiplet have $j_{1}=0$, so we need not worry about the $j_{1}=\pm 1/2$ contributions.
In the case of $I^{\text{(trifund)}}_{111}$, a generic operator that arises from the state-operator correspondence now has the form:
$\sum c^{n_{1}n_{2}}\partial^{n_{1}}_{+\dot{+}}\partial^{n_{2}}_{-\dot{+}}\Phi,$
where $\Phi$ denotes a trifundamental field. For clarity, we suppress the SU$(2)$ trifundamental indices on the field $\Phi$, which otherwise would appear as in equation (58). The $r$ and the $s$ constraints give
$c^{n_{1}n_{2}}=e^{i\pi(n_{1}-n_{2})/2}c^{n_{1}n_{2}},\qquad c^{n_{1}n_{2}}=(-1)^{n_{1}}c^{n_{2}n_{1}}.$ (79)
@@ B.2 The contribution $I^{\text{(trifund)}}$ of the half-hypermultiplet
$n_{1}-n_{2}\equiv 0\pmod{4},$
and the second constraint of (79) shows that the coefficients $c^{n_{1}n_{2}}$ for $n_{1}<n_{2}$ are determined in terms of the coefficients $c^{n_{1}n_{2}}$ for $n_{1}>n_{2}$, and for $n_{1}=n_{2}$ only the coefficients with even $n_{1}$ can be nonzero. Overall, we find that the number of independent coefficients $c^{n_{1}n_{2}}$ with a given $n_{1}+n_{2}$ is the coefficient of $t^{3(n_{1}+n_{2})}$ in
$\frac{1-t^{6}+t^{12}}{(1-t^{6})(1-t^{12})}\,.$
Altogether, the trifundamental index is
$I^{\text{(trifund)}}_{111}=8\frac{(1-t^{6}+t^{12})(t^{2}-t^{4})}{(1-t^{6})(1-t^{12})}.$ (80)
where the prefactor 8 comes from $2^{3}$ choices of the SU$(2)$ trifundamental indices, and the factor $(t^{2}-t^{4})$ comes from the contribution of the fields $q^{1}$ and $\overline{\psi}_{\dot{+}}$ as listed in Table 3, with $v=1$.
In the case of $I^{\text{(trifund)}}_{112}$, the constraints are
$e^{i\pi(n_{1}-n_{2})/2}c^{n_{1}n_{2}}=c^{n_{1}n_{2}},\qquad(-1)^{n_{1}}c^{n_{2}n_{1}}=-c^{n_{1}n_{2}},$ (81)
and the number of pairs $(n_{1},n_{2})$ with given $n_{1}+n_{2}$ is given by the coefficient of $t^{3(n_{1}+n_{2})}$ in the expansion of
$\frac{t^{6}}{(1-t^{6})(1-t^{12})}\,,$
@@
$I^{\text{(trifund)}}_{112}=8\frac{t^{6}(t^{2}-t^{4})}{(1-t^{6})(1-t^{12})}.$ (82)
In the case of $I^{\text{(trifund)}}_{113}$, the constraints are
$e^{i\pi(n_{1}-n_{2})/2}c^{n_{1}n_{2}}=-c^{n_{1}n_{2}},\qquad(-1)^{n_{1}}c^{n_{2}n_{1}}=c^{n_{1}n_{2}}.$ (83)
@@
$n_{1}-n_{2}\equiv 2\pmod{4},$
and the second one tells us to count all such $n_{1},n_{2}$ pairs with $n_{1}>n_{2}$. Their number, for a given $n_{1}+n_{2}$, is given by the coefficient of $t^{3(n_{1}+n_{2})}$ in the expansion of
$\frac{t^{6}}{(1-t^{6})(1-t^{12})},$
as we found in the computation of $I^{\text{(trifund)}}_{112}$ above. This gives
$I^{trifund}_{113}=8\frac{t^{6}(t^{2}-t^{4})}{(1-t^{6})(1-t^{12})}.$ (84)
In the case of $I^{\text{(trifund)}}_{114}$, the constraints are
$e^{i\pi(n_{1}-n_{2})/2}c^{n_{1}n_{2}}=-c^{n_{1}n_{2}},\qquad(-1)^{n_{1}}c^{n_{2}n_{1}}=-c^{n_{1}n_{2}}.$
Comparing the (114) constraints with the (113) constraints, we see that they have the same dimension of the space of solutions, so
$I^{\text{(trifund)}}_{114}=I^{\text{(trifund)}}_{113}=8\frac{t^{6}(t^{2}-t^{4})}{(1-t^{6})(1-t^{12})}.$ (85)
@@
Now, we study the index when one, two, or three of the holonomies in the triplet are the 5th holonomy. First, consider the case where there is only one non-abelian holonomy. Similarly to the argument we had previously, there are only 4 independent cases to consider: (115), (125), (135), (145).
In the case of $I^{\text{(trifund)}}_{115}$, the SU$(2)$ index of the third component of the trifundamental field matters, because the nonabelian holonomy does not act diagonally anymore. This consideration leads to the most general form of the operator:
$\sum c^{\mu n_{1}n_{2}}\partial^{n_{1}}_{+\dot{+}}\partial^{n_{2}}_{-\dot{+}}\Phi_{\mu},$
where $\mu=\pm 1$ is the gauge index of the trifundamental field with respect to the third SU$(2)$ factor. For clarity, we suppress the gauge indices of the first two SU$(2)$ factors. The constraints are now
$\sum e^{i\pi(n_{1}-n_{2})/2}c^{\mu n_{1}n_{2}}\partial^{n_{1}}_{+\dot{+}}\partial^{n_{2}}_{-\dot{+}}\Phi_{\mu}$ $=\sum e^{\mu\pi i/2}c^{\mu n_{1}n_{2}}\partial^{n_{1}}_{+\dot{+}}\partial^{n_{2}}_{-\dot{+}}\Phi_{\mu},$
$\sum(-1)^{n_{2}}c^{\mu n_{1}n_{2}}\partial^{n_{2}}_{+\dot{+}}\partial^{n_{1}}_{-\dot{+}}\Phi_{\mu}$ $=\sum(-1)^{(\mu-1)/2}c^{\mu n_{1}n_{2}}\partial^{n_{1}}_{+\dot{+}}\partial^{n_{2}}_{-\dot{+}}\Phi_{-\mu},$
@@
$e^{i\pi(n_{1}-n_{2})/2}c^{\mu n_{1}n_{2}}=e^{\mu\pi i/2}c^{\mu n_{1}n_{2}},\qquad(-1)^{n_{1}}c^{-\mu,n_{2}n_{1}}=(-1)^{(\mu-1)/2}c^{\mu n_{1}n_{2}}.$ (86)
For $\mu=1$, the first constraint of (86) implies
$n_{1}-n_{2}\equiv 1\pmod{4},$
and the second constraint instructs us to count all such $n_{1},n_{2}$ pairs with $\mu=1$, since the $\mu=-1$ coefficients are determined by the $\mu=1$ ones. For given $n_{1}+n_{2}$, this number is the coefficient of $t^{3(n_{1}+n_{2})}$ in the expansion of
$\frac{t^{3}+t^{9}}{(1-t^{6})(1-t^{12})},$
and the contribution to the index is
$I^{\text{(trifund)}}_{115}=4\frac{(t^{3}+t^{9})(t^{2}-t^{4})}{(1-t^{6})(1-t^{12})},$ (87)
where the $4$ comes from the $2^{2}$ choices of the first two SU$(2)$ trifundamental indices. The factor $t^{2}-t^{4}$ comes from the contribution of the half-hypermultiplet base fields $q^{1}$ and $\bar{\psi}_{\dot{+}}$.
In the case of $I^{\text{(trifund)}}_{125}$, the constraints are
$\sum e^{i\pi(n_{1}-n_{2})/2}c^{\mu n_{1}n_{2}}\partial^{n_{1}}_{+\dot{+}}\partial^{n_{2}}_{-\dot{+}}\Phi_{\mu}$ $=\sum e^{\mu\pi i/2}c^{\mu n_{1}n_{2}}\partial^{n_{1}}_{+\dot{+}}\partial^{n_{2}}_{-\dot{+}}\Phi_{\mu},$
$\sum(-1)^{n_{2}}c^{\mu n_{1}n_{2}}\partial^{n_{2}}_{+\dot{+}}\partial^{n_{1}}_{-\dot{+}}\Phi_{\mu}$ $=-\sum(-1)^{(\mu-1)/2}c^{\mu n_{1}n_{2}}\partial^{n_{1}}_{+\dot{+}}\partial^{n_{2}}_{-\dot{+}}\Phi_{-\mu}.$
The only difference from the previous $115$ case is the minus sign for the $s$ condition, but this sign does not change the count, and
$I^{\text{(trifund)}}_{125}=I^{\text{(trifund)}}_{115}=4\frac{(t^{3}+t^{9})(t^{2}-t^{4})}{(1-t^{6})(1-t^{12})}.$ (88)
In the case of $I^{\text{(trifund)}}_{135}$, the constraints are
$\sum e^{i\pi(n_{1}-n_{2})/2}c^{\mu n_{1}n_{2}}\partial^{n_{1}}_{+\dot{+}}\partial^{n_{2}}_{-\dot{+}}\Phi_{\mu}$ $=-\sum e^{\mu\pi i/2}c^{\mu n_{1}n_{2}}\partial^{n_{1}}_{+\dot{+}}\partial^{n_{2}}_{-\dot{+}}\Phi_{\mu},$
$\sum(-1)^{n_{2}}c^{\mu n_{1}n_{2}}\partial^{n_{2}}_{+\dot{+}}\partial^{n_{1}}_{-\dot{+}}\Phi_{\mu}$ $=\sum(-1)^{(\mu-1)/2}c^{\mu n_{1}n_{2}}\partial^{n_{1}}_{+\dot{+}}\partial^{n_{2}}_{-\dot{+}}\Phi_{-\mu}.$
@@
$I^{\text{(trifund)}}_{135}=I^{\text{(trifund)}}_{115}=4\frac{(t^{3}+t^{9})(t^{2}-t^{4})}{(1-t^{6})(1-t^{12})}.$ (89)
@@
$I^{\text{(trifund)}}_{145}=I^{\text{(trifund)}}_{115}=4\frac{(t^{3}+t^{9})(t^{2}-t^{4})}{(1-t^{6})(1-t^{12})}.$ (90)
@@
This completes the computation for the cases where only one of the three holonomies is nonabelian. We now consider the four cases where two of the holonomies are nonabelian. Those are: (155), (255), (355), and (455).
@@
In the case of $I^{trifund}_{155}$, the second and the third SU$(2)$ trifundamental gauge indices matter, since the holonomies will act non-diagonally on them separately. Suppressing only the gauge indices of the first SU$(2)$ factor, we write the components of the trifundamental field as $\Phi_{\mu\nu}$ ($\mu,\nu=\pm 1$), and the general operator combination is written as
$\sum c^{\mu\nu n_{1}n_{2}}\partial^{n_{1}}_{+\dot{+}}\partial^{n_{2}}_{-\dot{+}}\Phi_{\mu\nu}.$
The constraints are
$\sum e^{i\pi(n_{1}-n_{2})/2}c^{\mu\nu n_{1}n_{2}}\partial^{n_{1}}_{+\dot{+}}\partial^{n_{2}}_{-\dot{+}}\Phi_{\mu\nu}$ $=\sum e^{i\pi(\mu+\nu)/2}c^{\mu\nu n_{1}n_{2}}\partial^{n_{1}}_{+\dot{+}}\partial^{n_{2}}_{-\dot{+}}\Phi_{\mu\nu},$
$\sum(-1)^{n_{2}}c^{\mu\nu n_{1}n_{2}}\partial^{n_{2}}_{+\dot{+}}\partial^{n_{1}}_{-\dot{+}}\Phi_{\mu\nu}$ $=\sum(-1)^{(\mu+\nu)/2-1}c^{\mu\nu n_{1}n_{2}}\partial^{n_{1}}_{+\dot{+}}\partial^{n_{2}}_{-\dot{+}}\Phi_{-\mu,-\nu},$
from which we obtain
$e^{i\pi(n_{1}-n_{2})/2}c^{\mu\nu n_{1}n_{2}}=e^{i\pi(\mu+\nu)/2}c^{\mu\nu n_{1}n_{2}},\quad(-1)^{n_{1}}c^{-\mu,-\nu,n_{2}n_{1}}=(-1)^{(\mu+\nu)/2-1}c^{\mu\nu n_{1}n_{2}}.$ (91)
We proceed by counting the number of linearly independent $c^{\mu\nu n_{1}n_{2}}$’s for each $(\mu,\nu)$ separately.
  • For $\mu=\nu=1$, the first equation of (91) implies
    $n_{1}-n_{2}\equiv 2\pmod{4},$ (92)
    and the second equation says that the coefficients with $\mu=\nu=-1$ are determined by those with $\mu=\nu=1$. The number of pairs satisfying (92) for a given $n_{1}+n_{2}$ is the coefficient of $t^{3(n_{1}+n_{2})}$ in the expansion of
    $W_{11}=\frac{2t^{6}}{(1-t^{6})(1-t^{12})}.$
  • For $\mu=1,\nu=-1$, the first constraint of (91) implies
    $n_{1}-n_{2}\equiv 0\pmod{4},$ (93)
    and the second constraint determines the coefficients with $\mu=-1$ and $\nu=1$ in terms of those with $\mu=1$ and $\nu=-1$. The generating function is
    $W_{1,-1}=\frac{1+t^{12}}{(1-t^{6})(1-t^{12})}.$
Adding these together, we have
$I^{\text{(trifund)}}_{155}=2(W_{11}+W_{1,-1})(t^{2}-t^{4})=\frac{2(1+2t^{6}+t^{12})(t^{2}-t^{4})}{(1-t^{6})(1-t^{12})},$ (94)